what is the magnitude of the electric field at a point midway between a-6.5uC and +5.4uC charge 8.0cm apart ? assume no other charge nearby . express your answer using two significant figures.


1
Expert's answer
2021-10-18T11:00:37-0400

"E = \frac{kQ_1}{r_1^2} + \frac{kQ_2}{r^2_2} \\

E = \frac{9 \times 10^9 \times (-6.5 \times 10^{-6})}{0.04^2} - \frac{9 \times 10^9 \times 5.4 \times 10^{-6}}{0.04^2} \\

= 6.7 \times 10^7 \;N/C"

In negative x direction